OpenStudy (anonymous):

The table below shows the number of marbles of different colors in a bag: Marble Experiment Color of Marbles Number of Marbles Red 1 White 5 Black 10 Ursula draws a marble from the bag randomly without looking. She then draws another marble from the bag without replacing the first one. Which expression shows the probability of drawing black marbles in both the trials? 10 over 16 added to 9 over 16 10 over 16 added to 9 over 15 10 over 16 multiplied by 9 over 16 10 over 16 multiplied by 9 over 15

OpenStudy (anonymous):

@knov

OpenStudy (kropot72):

\[\large P(2\ black)=\frac{10C2}{16C2}\]

OpenStudy (anonymous):

so its d? 10 over 16 multiplied by 9 over 15 @kropot72

OpenStudy (anonymous):

@ILovePuppiesLol

OpenStudy (anonymous):

@ILovePuppiesLol

ILovePuppiesLol (ilovepuppieslol):

i found it

ILovePuppiesLol (ilovepuppieslol):

@rebeccaxhawaii help

OpenStudy (anonymous):

@ILovePuppiesLol pleaseee helppp meee

OpenStudy (kropot72):

\[\large \frac{10C2}{16C2}=\frac{10!}{2!8!}\times\frac{2!14!}{16!}=\frac{14\times13\times12\times11\times10\times9}{16\times15\times14\times13\times12\times11}\]

OpenStudy (kropot72):

@ap1usstudent Now do some cancellation to confirm the correct answer.